If x= 0 is rational and y irrational, prove that x + y,x − y,xy,x/y and y/x are all irrational.


1
Expert's answer
2021-01-19T04:21:50-0500

Here we have "(x\neq0)\in\mathbb{Q}" and y"\in\mathbb{Q}^c"


  • "x+y"

Here let's assume this is rational.

Now, lets take x=3 and y="\sqrt{2}"

So, "x+y=3+\sqrt{2}" which is irrational.

So, we have a counter-example to show that "x+y" is irrational. Hence our claim is false and is irrational.


  • "x-y"

Here let's assume this is rational.

Now, lets take x=3 and y="\sqrt{2}"

So, "x-y=3-\sqrt{2}" which is irrational.

So, we have a counter-example to show that "x-y" is irrational. Hence our claim is false and is irrational.


  • "xy"

Here let's assume this is rational.

Now, lets take x=3 and y="\sqrt{2}"

So, "xy=3\sqrt{2}" which is irrational.

So, we have a counter-example to show that "xy" is irrational. Hence our claim is false and is irrational.


  • "\frac{x}{y}"

Here let's assume this is rational.

Now, lets take x=3 and y="\sqrt{2}"

So, "\frac{x}{y}=\frac{3}{\sqrt{2}}" which is irrational.

So, we have a counter-example to show that "\frac{x}{y}" is irrational. Hence our claim is false and is irrational.


  • "\frac{y}{x}"

Here let's assume this is rational.

Now, lets take x=3 and y="\sqrt{2}"

So, "\frac{y}{x}=\frac{\sqrt{2}}{3}" which is irrational.

So, we have a counter-example to show that "\frac{y}{x}" is irrational. Hence our claim is false and is irrational.